STRUCTRURAL DETERMINATION OF SILACYCLOBUTANE AND SILACYCLOPENTANE USING FOURIER TRANSFORM MICROWAVE (FTMW) AND CHIRPED PULSE FOURIER TRANSFORM MI-CROWAVE (cp-FTMW) SPECTROSCOPY

\underline{ZIQIU CHEN}, CODY VAN DIJK AND JENNIFER VAN WIJNGAARDEN, Department of Chemistry, University of Manitoba, Winnipeg MB R3T 2N2 Canada.

The pure rotational spectra of the ground states of silacyclobutane (SCB) and silacyclopentane (SCP) were measured in a supersonic jet in the 6-24 GHz range using Fourier transform microwave spectroscopy and the chirped-pulse variant of this technique. Heavy atom isotopic substitution for the silicon and each of the carbon atoms within the rings enabled the accurate determination of the r_{s} and r_{0} structural parameters of the ring backbones of both SCB and SCP. For SCB, splitting due to ring inversion in the ground state has been observed and analyzed.